13. What is the molarity of a 2.66 L aqueous solution containing 56.9 g of NaBr?

14. If 3.08 g of Ca(OH)2 is dissolved in enough water to make 0.875 L of solution, what is the molarity of the Ca(OH)2?

15. What mass of HCl is present in 825 mL of a 1.25 M solution?

16. What mass of isopropyl alcohol (C3H8O) is dissolved in 2.050 L of a 4.45 M aqueous C3H8O solution?
Answer #1

13)

Molar mass of NaBr,

MM = 1*MM(Na) + 1*MM(Br)

= 1*22.99 + 1*79.9

= 102.89 g/mol

mass(NaBr)= 56.9 g

use:

number of mol of NaBr,

n = mass of NaBr/molar mass of NaBr

=(56.9 g)/(1.029*10^2 g/mol)

= 0.553 mol

volume , V = 2.66 L

use:

Molarity,

M = number of mol / volume in L

= 0.553/2.66

= 0.2079 M

Answer: 0.208 M

14)

Molar mass of Ca(OH)2,

MM = 1*MM(Ca) + 2*MM(O) + 2*MM(H)

= 1*40.08 + 2*16.0 + 2*1.008

= 74.096 g/mol

mass(Ca(OH)2)= 3.08 g

use:

number of mol of Ca(OH)2,

n = mass of Ca(OH)2/molar mass of Ca(OH)2

=(3.08 g)/(74.1 g/mol)

= 4.157*10^-2 mol

volume , V = 0.875 L

use:

Molarity,

M = number of mol / volume in L

= 4.157*10^-2/0.875

= 4.751*10^-2 M

Answer: 4.75*10^-2 M

Only 1 question at a time please